If f(x)=a/(3-x) where X doesn't equal to 3 and g(x)=11-bx^2, f^(5)=4/5 and fg(2)=1/2
Find the values of a and b!

Grade:10

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

F(5)= 4/5 = a/(-2)

a = -8/5

Similarly use last equation and value of a to obtain b, where the function is f(g(x)) ie
f(g(2))

Regards
Yashraj
3 years ago
F(5)=4/5                                                                                                                                                a=-8/5
g(2)=11-4b then
put value of g(2)in last equation
then get value of b
 
with regards
these types of questions are simple practice